How do you factor the equation 2a^2 + 8a = 0? Thank you.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

How do you factor the equation 2a^2 + 8a = 0?
-----------
You don't.
Factor 2a^2 + 8a
Hint:
Both coefficients are even.
Both terms have an a.

Answer by stanbon(75887)

factor the equation 2a^2 + 8a = 0? Thank you.
----
Common Factor: 2a
---
2a(a+4) = 0
Either 2a is 0 or a+4 = 0
a = 0 or a = -4
==============================
cheers,
Stan H.
===========